OpenStudy (anonymous):

A family refers to a group of two or more people related by birth marriage, or adoption who reside together. In 2000, in Country A, the average family net worth was $350,000, and there about 7.6*10^7 families. Calculate the total family net worth in Country A in 2000. County A's total family net worth is?

OpenStudy (anonymous):

3.5*10^5*7.6*10^7=26.6*10^12=2.66*10^13